# Direct Data-Driven Predictive Control for a Three-dimensional Cable-Driven Soft Robotic Arm

### v1

@@ Abstract
Soft robots offer significant advantages in safety and adaptability, yet achieving precise and dynamic control remains a major challenge due to their inherently complex and nonlinear dynamics. Recently, Data-enabled Predictive Control (DeePC) has emerged as a promising model-free approach that bypasses explicit system identification by directly leveraging input–output data. While DeePC has shown success in other domains, its application to soft robots remains underexplored, particularly for three-dimensional (3D) soft robotic systems. This paper addresses this gap by developing and experimentally validating an effective DeePC framework on a 3D, cable-driven soft arm. Specifically, we design and fabricate a soft robotic arm with a thick tubing backbone for stability, a dense silicone body with large cavities for strength and flexibility, and rigid endcaps for secure termination. Using this platform, we implement DeePC with singular value decomposition (SVD)-based dimension reduction for two key control tasks: fixed-point regulation and trajectory tracking in 3D space. Comparative experiments with a baseline model-based controller demonstrate DeePC’s superior accuracy, robustness, and adaptability, highlighting its potential as a practical solution for dynamic control of soft robots.
@@ I INTRODUCTION
However, this promise comes with profound control challenges. The very sources of soft robots’ advantages, deformable materials and continuum structures, also lead to extreme complexity in modeling and control [8]. A primary difficulty lies in their complex nonlinear dynamics: materials such as silicone exhibit nonlinear stress–strain behavior, and under pneumatic, electrical, or tendon actuation, their responses often involve strong nonlinearity, hysteresis, and creep [9]. Another challenge is the infinite-dimensional state space. Unlike rigid robots, whose configurations can be described by a finite set of joint angles and velocities, a continuously deformable body theoretically requires specifying the displacement of every point, resulting in an infinite-dimensional system. Collectively, these factors make soft robot control particularly challenging [10].
@@
To address the complex dynamics of soft robots, researchers have proposed numerous model-based control strategies; however, these methods often struggle to balance accuracy and computational complexity [9]. For instance, the Piecewise Constant Curvature (PCC) model is the most widely used simplification [11]. It assumes each section bends as a circular arc with constant curvature, reducing the configuration to only a few parameters. Yet, PCC’s validity depends on this assumption, which is frequently violated in practice due to gravity, environmental interactions, and non-uniform actuation or materials, resulting in significant model–plant mismatch and degraded control performance [12]. The finite element method (FEM) offers a more accurate alternative, capable of capturing complex geometries, nonlinear material properties, and contact interactions between soft robots and their environments. However, this accuracy comes at the cost of high computational demands [13]. To address this, Navez et al. proposed a condensed FEM model that combines the efficiency of machine learning with the physical interpretability of FEM-based mechanics [14]. While promising, it relies on quasi-static assumptions, requires extensive FEM-generated training data, and shows limited generalization to highly nonlinear or dynamic scenarios.
@@
Given the limitations of first-principles modeling, data-driven methods provide a promising alternative for soft robot control. Instead of deriving exact physical models, these approaches learn input–output mappings or latent dynamics directly from experimental data. Reinforcement learning (RL) offers an appealing model-free control strategy for systems that are difficult to model; however, its use is hindered by low sample efficiency, the simulation-to-reality gap, and safety concerns associated with exploration [15, 16]. Koopman operator theory provides a distinct, data-driven approach. Unlike RL, it is not model-free; rather, it identifies an explicit linear model from data that represents the dynamics of the underlying nonlinear system. Bruder et al. [17] demonstrated a Koopman-based method that enabled real-time trajectory tracking of soft robots with high computational efficiency. Nevertheless, its effectiveness depends heavily on the choice of basis functions, and for more complex systems, the model dimension can grow rapidly and become computationally intractable.
@@
In contrast to RL’s fully model-free nature and Koopman operator theory’s explicit modeling, data-enabled predictive control (DeePC) adopts a nonparametric, model-implicit approach by leveraging the Hankel matrix of past input–output data as its predictive model [18]. DeePC has been successfully applied in several domains, including large-scale damping control in power systems [19], building automation [20], quadrotor trajectory tracking [21], and power-converter control [22]. Its robustness, ability to handle nonlinearity and uncertainty, and nonparametric nature align closely with the challenges of soft robot control [23], yet related applications remain at an early stage. Wang et al. were the first to apply DeePC to soft robots [24], but their validation was limited to a planar, pneumatically actuated system restricted to one plane of motion. Extending DeePC to three-dimensional soft arms and validating it in high-precision 3D dynamic tasks thus remains a key open problem. To address this gap, we build on our previous work on a modular cable-driven soft robotic arm, for which we developed a nonlinear kinematic model and demonstrated open-loop control [25]. This study advances that foundation by (i) optimizing the robotic arm design, (ii) introducing, for the first time, a DeePC framework with singular value decomposition (SVD)-based dimension reduction, systematically implemented and experimentally validated on the developed soft robotic arm platform for two control tasks: fixed-point regulation and dynamic trajectory tracking in 3D space, and (iii) open-sourcing the arm designs and control codes to facilitate future research [^1] Codes and designs: https://github.com/chengoy30/DeePC-for-Soft-Robot-Control.
@@ II-A Robot design and fabrication
For applying DeePC, we initially tested our previous robot design in [25]. However, through repeated testing and several design iterations, we found that practical applications require greater stiffness and load-bearing capacity while still preserving the inherent flexibility of soft robots. To address this, the final design employed a thicker flexible backbone and denser silicone, providing the necessary structural strength while maintaining compliance. The resulting soft robot consists of two end caps, a flexible backbone, and three helix-like spiral structures that incorporate thin strings for cable routing. The detailed structure and sequential fabrication process are shown in Fig. 11, while the final cured soft robot is presented in Fig. 1. The end caps (Fig. 1) were 3D-printed using PETG, with anchor points designed to prevent detachment of the silicone after casting, and further shaped to support the mounting and stacking of multiple arm segments. A flexible backbone was formed using clear Masterkleer soft PVC tubing (McMaster-Carr) with an outer diameter of 8 mm and an inner diameter of 3 mm. The molds for silicone casting (Fig. 1) were also fabricated via 3D printing, and Ecoflex 00-50 silicone was used as the elastomer material. High-strength Kevlar threads served dual purposes: as actuation cables and as reinforcement in the helix-like spiral structures.
@@
The fabrication process begins by aligning a 88 mm length of flexible PVC backbone tubing along the central axis, with both ends secured to the 3D-printed end caps. A helix-like spiral structure is then formed by winding Kevlar thread around 2 mm diameter metal rods, with the helix endpoints anchored to the end caps. To create internal cavities, three 3D-printed beams were positioned around the backbone at 120° intervals. Three 3D-printed shells were subsequently used to encapsulate the end caps, backbone, helix structure, and cavity beams. One end of the mold was sealed with a 3D-printed cap, and petroleum jelly was applied at potential leakage points. Ecoflex 00-50 silicone was then poured through openings in the top end cap and left to cure for 4 hours. After curing, the rods and beams were removed, and the mold shell along with the bottom cap was disassembled, yielding the completed soft robot. The final dimensions of the robot are 29 mm in diameter and 90 mm in length.
For actuation, Kevlar threads were routed through the 2 mm helix structures embedded within the robot. The helix structures serve to protect the surrounding silicone from damage caused by repeated pulling and stretching of these tendon-like cables. One end of each thread was tied to a metal stopper and anchored to an end cap, while the other end exited the robot and connected to stepper motors via a 3D-printed pulley system.
II-B Control board
@@ III-A Data-enabled Predictive Control (DeePC)
We begin with an overview of DeePC [18]. By linearizing the dynamics around a nominal operating point, the behavior of a soft robot can be approximated by a discrete-time linear time-invariant (LTI) system:
@@
The core idea of DeePC is to bypass explicit system identification by directly utilizing offline input–output data rather than a parametric model [18]. This approach is grounded in behavioral systems theory, which represents the system’s dynamics using a block Hankel matrix [26]. For the specific case of our soft robotic arm, we begin by conducting an offline data collection experiment. A predefined sequence of control inputs $u^{\mathrm{d}}=\begin{bmatrix}u^{\mathrm{d}}(0)^{\top},u^{\mathrm{d}}(1)^{\top},\ldots,u^{\mathrm{d}}(T-1)^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{m\times T}$, which denotes the incremental lengths of the three actuation cables, is applied to the robotic arm. Simultaneously, the resulting trajectory of the end-effector is recorded as the output sequence $y^{\mathrm{d}}=\begin{bmatrix}y^{\mathrm{d}}(0)^{\top},y^{\mathrm{d}}(1)^{\top},\ldots,y^{\mathrm{d}}(T-1)^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{p\times T}$. This collected input-output dataset is then structured into a block Hankel matrix of depth $L\in\mathbb{Z}$. This data matrix serves as a non-parametric model of the robot’s dynamics and is defined as:
@@
A prerequisite for invoking the lemma is that the input sequence used for offline data collection must be persistently exciting (PE). An input $u^{\mathrm{d}}$ is said to be PE of order $L$ if its corresponding Hankel matrix $\mathcal{H}_{L}(u^{\mathrm{d}})$ has full row rank. To satisfy this condition in practice, a rich variety of input signals was applied to the soft robotic arm during the offline phase, ensuring that the arm explored a comprehensive portion of its workspace and that the resulting data matrix fully captured its dynamics.
According to the lemma, for any controllable soft robotic arm system driven by a PE input $u^{\mathrm{d}}$, any feasible motion trajectory of length $L$ can be accurately represented, and any future motion that the robot can achieve can be regarded as a linear combination of the behaviors it has exhibited in the past. Building on this principle, DeePC partitions the Hankel matrices into past and future components, which are then used to formulate a constrained optimal control problem. Specifically, let $T_{\mathrm{ini}}$, $N\in\mathbb{Z}$, and $L=T_{\mathrm{ini}}+N$. The following block matrices are defined:
$\begin{bmatrix}U_{p}\\ U_{f}\end{bmatrix}:=\mathcal{H}_{L}\!\left(u^{\mathrm{d}}\right),\quad\begin{bmatrix}Y_{p}\\ Y_{f}\end{bmatrix}=\mathcal{H}_{L}\!\left(y^{\mathrm{d}}\right),$ (4)
@@
where $y_{r}$ is the reference trajectory, $Q$, $R$ are weighting matrices, $u_{\mathrm{ini}}=\begin{bmatrix}u(t-T_{\mathrm{ini}})^{\top},u(t-T_{\mathrm{ini}}+1)^{\top},\ldots,u(t-1)^{\top}\end{bmatrix}^{\top}$ and $y_{\mathrm{ini}}=\begin{bmatrix}y(t-T_{\mathrm{ini}})^{\top},y(t-T_{\mathrm{ini}}+1)^{\top},\ldots,y(t-1)^{\top}\end{bmatrix}^{\top}$ are the input and output sequences within a past time horizon of length $T_{\mathrm{ini}}$, $u=\begin{bmatrix}u(t)^{\top},u(t+1)^{\top},\ldots,u(t+N-1)^{\top}\end{bmatrix}^{\top}$ and $y=\begin{bmatrix}y(t)^{\top},y(t+1)^{\top},\ldots,y(t+N-1)^{\top}\end{bmatrix}^{\top}$ are the input and output sequences within a prediction horizon of length $N$, and $\mathcal{U}$, $\mathcal{Y}$ represent the admissible sets of inputs and outputs, respectively.
@@ III-B Robust DeePC via Regularization
Although DeePC and MPC are equivalent under ideal conditions (i.e., a controllable LTI system with a persistently exciting input) [18], this equivalence rarely holds for the real-world soft robotic arm. The ideal assumptions are challenged by multiple factors: (i) measurement noise from sensors and external disturbances (e.g., airflow, vibrations), and (ii) the highly nonlinear dynamics inherent to soft materials (e.g., viscoelasticity, hysteresis), which cannot be fully captured by a finite dataset. These imperfections mean that a newly observed trajectory $(u_{ini},y_{ini})$ may not be perfectly linearly represented in the historical database $(U_{p},Y_{p})$. This mismatch makes the equality constraint in the optimization problem impossible to satisfy, leading to infeasibility. To restore feasibility and improve robustness against these real-world effects, the standard DeePC formulation is often augmented with regularization. Common approaches include:
@@
By truncating this decomposition to the top $r$ singular values, which capture the system’s principal dynamics, a new condensed data matrix $\tilde{\mathcal{H}}_{L}\in\mathbb{R}^{q_{1}\times r}$ is obtained:
@@
$\min_{\bar{g},u,y,\sigma_{y}}$ $\|y-y_{r}\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}+\lambda_{g}\|\tilde{g}\|_{2}^{2}$ (9)
@@
This formulation provides a significant computational advantage: the dimension of the optimization variable is reduced from the dataset length (TL+1)(T-L+1) to the approximation rank $r$. Such reduction greatly improves the efficiency of solving the DeePC problem, thereby making it more suitable for real-time control of the soft robotic arm [24]. The choice of $r$ can be guided by inspecting the decay of the singular values or determined empirically based on control performance.

### v2

@@ Abstract
Soft robots offer significant advantages in safety and adaptability, yet achieving precise and dynamic control remains a major challenge due to their inherently complex and nonlinear dynamics. Recently, Data-enabled Predictive Control (DeePC) has emerged as a promising model-free method that bypasses explicit system identification by directly leveraging input–output data. While DeePC has shown success in other domains, its application to soft robots remains underexplored, particularly for 3D soft robotic systems. This paper addresses this gap by developing and experimentally validating an effective DeePC framework on a 3D cable-driven soft arm. Specifically, we design and fabricate a soft robotic arm with a thick tubing backbone for stability, a dense silicone body with large cavities for strength and flexibility, and rigid endcaps for secure termination. Using this platform, we implement DeePC with singular value decomposition (SVD)-based dimension reduction for two key control tasks: fixed-point regulation and trajectory tracking in 3D space. Comparative experiments with a baseline model-based controller demonstrate DeePC’s superior accuracy and adaptability, highlighting its potential as a practical solution for dynamic control of soft robots.
@@ I INTRODUCTION
However, this promise comes with profound control challenges. The very sources of soft robots’ advantages, deformable materials and continuum structures, also lead to extreme complexity in modeling and control [24]. A primary difficulty lies in their highly nonlinear dynamics: materials such as silicone exhibit nonlinear stress–strain behavior, and under pneumatic, electrical, or tendon actuation, their responses often involve strong nonlinearity, hysteresis, and creep [9]. Another challenge is the infinite-dimensional state space. Unlike rigid robots, whose configurations can be described by a finite set of joint angles and velocities, a continuously deformable body theoretically requires specifying the displacement of every point, resulting in an infinite-dimensional system. Collectively, these factors make soft robot control particularly challenging [28].
@@
To address the complex dynamics of soft robots, researchers have proposed numerous model-based control strategies; however, these methods often struggle to balance accuracy and computational complexity [9]. For instance, the Piecewise Constant Curvature (PCC) model is the most widely used simplification [16]. It assumes each section bends as a circular arc with constant curvature, reducing the configuration to only a few parameters. Yet, PCC’s validity depends on this assumption, which is frequently violated in practice due to gravity, environmental interactions, and non-uniform actuation or materials, resulting in significant model–plant mismatch and degraded performance [21]. The finite element method (FEM) offers a more accurate alternative, capable of capturing complex geometries, nonlinear material properties, and contact interactions between soft robots and their environments. However, this accuracy comes at the cost of high computational demands [4]. To address this, Navez et al. proposed a condensed FEM model that combines the efficiency of machine learning with the physical interpretability of FEM-based mechanics [18]. While promising, it relies on quasi-static assumptions, requires extensive FEM-generated training data, and shows limited generalization to highly nonlinear or dynamic scenarios.
@@
Given the limitations of first-principles modeling, data-driven methods provide a promising alternative for soft robot control. Instead of deriving exact physical models, these approaches learn input–output mappings or latent dynamics directly from experimental data. Reinforcement learning (RL) offers an appealing model-free control strategy for systems that are difficult to model; however, its use is hindered by low sample efficiency, the simulation-to-reality gap, and safety concerns associated with exploration [19, 11]. Koopman operator theory provides a distinct, data-driven approach. Unlike RL, it is not model-free; rather, it identifies an explicit linear model from data that represents the dynamics of the underlying nonlinear system. Bruder et al. [2] demonstrated a Koopman-based method that enabled real-time trajectory tracking of soft robots with high computational efficiency. However, its effectiveness relies heavily on the choice of basis functions, and for more complex systems, the model dimension can become computationally intractable.
@@
In contrast to RL’s fully model-free nature and Koopman operator theory’s explicit modeling, data-enabled predictive control (DeePC) adopts a nonparametric, model-implicit approach by leveraging the Hankel matrix of past input–output data as its predictive model [7]. DeePC has been successfully applied in several domains, including large-scale damping control in power systems [13], building automation [5], quadrotor trajectory tracking [10], and power-converter control [14]. Its robustness, ability to handle nonlinearity and uncertainty, and nonparametric nature align closely with the challenges of soft robot control [15], yet related applications remain at an early stage. Wang et al. were the first to apply DeePC to soft robots [23], but their validation was limited to a planar, pneumatically actuated system restricted to one plane of motion. Extending DeePC to three-dimensional soft arms and validating it in high-precision 3D dynamic tasks thus remains a key open problem. To address this gap, we build on our previous work on a modular cable-driven soft robotic arm, for which we developed a nonlinear kinematic model and demonstrated open-loop control [20]. This study advances that foundation by (i) optimizing the robotic arm design, (ii) introducing a DeePC framework with singular value decomposition (SVD)-based dimension reduction, systematically implemented and experimentally validated on the developed soft robotic arm platform for two control tasks: fixed-point regulation and dynamic trajectory tracking in 3D space, and (iii) open-sourcing the arm designs and control codes to facilitate future research [^1] Codes and designs: https://github.com/chengoy30/DeePC-for-Soft-Robot-Control.
@@ II-A Robot design and fabrication
To apply DeePC, we initially tested our previous robot design in [20]. However, through repeated testing and several design iterations, we found that practical applications require greater stiffness and load-bearing capacity while preserving the inherent flexibility of soft robots. To address this, the final design employed a thicker flexible backbone and denser silicone, providing the necessary structural strength while maintaining compliance. The resulting soft robotic arm consists of two end caps, a flexible backbone, and three helix-like spiral structures that incorporate Kevlar threads for cable routing. The detailed structure and sequential fabrication process are shown in Fig. 11, while the final cured soft robot is presented in Fig. 1. The end caps (Fig. 1) were 3D-printed using PETG, with anchor points designed to prevent detachment of the silicone after casting, and further shaped to support the mounting and stacking of multiple arm segments. A flexible backbone was formed using clear Masterkleer soft PVC tubing (McMaster-Carr) with an outer diameter of 8 mm and an inner diameter of 3 mm. The molds for silicone casting (Fig. 1) were also fabricated via 3D printing, and Ecoflex 00-50 silicone was used as the elastomer material. High-strength Kevlar threads served dual purposes: as actuation cables and as reinforcement in the helix-like spiral structures.
@@
II-B Actuation and control system
@@ III-A Data-enabled Predictive Control (DeePC)
We begin with an overview of DeePC [7]. By linearizing the system dynamics around a nominal operating point, the behavior of a soft robot can be approximated by a discrete-time linear time-invariant (LTI) system:
@@
The core idea of DeePC is to bypass explicit system identification by directly utilizing offline input–output data rather than a parametric model [7]. This approach is grounded in behavioral systems theory, which represents the system’s dynamics using a block Hankel matrix [26]. For the specific case of our soft robotic arm, we begin by conducting an offline data collection experiment. A predefined sequence of control inputs $u^{\mathrm{d}}=\begin{bmatrix}u^{\mathrm{d}}(0)^{\top},u^{\mathrm{d}}(1)^{\top},\ldots,u^{\mathrm{d}}(T-1)^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{mT}$, which denotes the incremental lengths of the three actuation cables, is applied to the robotic arm. Simultaneously, the resulting trajectory of the end-effector is recorded as the output sequence $y^{\mathrm{d}}=\begin{bmatrix}y^{\mathrm{d}}(0)^{\top},y^{\mathrm{d}}(1)^{\top},\ldots,y^{\mathrm{d}}(T-1)^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{pT}$. This collected input-output dataset is then structured into a block Hankel matrix of depth $L\in\mathbb{Z}$. This data matrix serves as a non-parametric model of the robot’s dynamics and is defined as:
@@
A prerequisite for invoking the lemma is that the input sequence used for offline data collection must be persistently exciting (PE). An input $u^{\mathrm{d}}$ is said to be PE of order $L$ if its corresponding Hankel matrix $\mathcal{H}_{L}(u^{\mathrm{d}})$ has full row rank. To satisfy this condition in practice, a rich variety of input signals were applied to the soft robotic arm during the offline phase, ensuring that the arm explored a comprehensive portion of its workspace and that the resulting data matrix fully captured its dynamics.
According to the lemma, for any controllable soft robotic arm system driven by a PE input $u^{\mathrm{d}}$, any feasible motion trajectory of length $L$ can be accurately represented, and any future motion that the robot can achieve can be represented as a linear combination of the behaviors it has exhibited in the past. Building on this principle, DeePC partitions the Hankel matrices into past and future components, which are then used to formulate a constrained optimal control problem. Specifically, let $T_{\mathrm{ini}}$, $N\in\mathbb{Z}$, and $L=T_{\mathrm{ini}}+N$. The following block matrices are defined:
$\begin{bmatrix}U_{p}\\ U_{f}\end{bmatrix}:=\mathcal{H}_{L}\!\left(u^{\mathrm{d}}\right),\quad\begin{bmatrix}Y_{p}\\ Y_{f}\end{bmatrix}:=\mathcal{H}_{L}\!\left(y^{\mathrm{d}}\right),$ (4)
@@
where $y_{r}$ is the reference trajectory, $Q$, $R$ are weighting matrices, $u_{\mathrm{ini}}=\begin{bmatrix}u(t-T_{\mathrm{ini}})^{\top},u(t-T_{\mathrm{ini}}+1)^{\top},\ldots,u(t-1)^{\top}\end{bmatrix}^{\top}$ and $y_{\mathrm{ini}}=\begin{bmatrix}y(t-T_{\mathrm{ini}})^{\top},y(t-T_{\mathrm{ini}}+1)^{\top},\ldots,y(t-1)^{\top}\end{bmatrix}^{\top}$ are the input and output sequences within a past time horizon of length $T_{\mathrm{ini}}$, $u=\begin{bmatrix}u(t)^{\top},u(t+1)^{\top},\ldots,u(t+N-1)^{\top}\end{bmatrix}^{\top}$ and $y=\begin{bmatrix}y(t)^{\top},y(t+1)^{\top},\ldots,y(t+N-1)^{\top}\end{bmatrix}^{\top}$ are the input and output sequences within a prediction horizon of length $N$, and $\mathcal{U}$, $\mathcal{Y}$ represent the admissible sets of inputs and outputs.
@@ III-B Robust DeePC via Regularization
Although DeePC and MPC are equivalent under ideal conditions (i.e., a controllable LTI system with a persistently exciting input) [7], this equivalence rarely holds for the real-world soft robotic arm. The ideal assumptions are challenged by multiple factors: (i) measurement noise from sensors and external disturbances (e.g., airflow, vibrations), and (ii) the highly nonlinear dynamics inherent to soft materials (e.g., viscoelasticity, hysteresis), which cannot be fully captured by a finite dataset. These imperfections mean that a newly observed trajectory $(u_{\mathrm{ini}},y_{\mathrm{ini}})$ may not be perfectly linearly represented in the historical database $(U_{p},Y_{p})$. This mismatch makes the equality constraint in the optimization problem impossible to satisfy, leading to infeasibility. To restore feasibility and improve robustness against these real-world effects, the standard DeePC formulation is often augmented with regularization. Common approaches include:
@@
By truncating this decomposition to the top $r$ singular values, which capture the system’s principal dynamics, a new condensed data matrix $\bar{\mathcal{H}}_{L}\in\mathbb{R}^{q_{1}\times r}$ is obtained:
@@
$\min_{\bar{g},u,y,\sigma_{y}}$ $\|y-y_{r}\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}+\lambda_{g}\|\bar{g}\|_{2}^{2}$ (9)
@@
This formulation provides a significant computational advantage: the dimension of the optimization variable is reduced from the dataset length (TL+1)(T-L+1) to the approximation rank $r$. Such a reduction makes the approach more suitable for real-time control of the soft robotic arm [23]. The choice of $r$ can be guided by inspecting the decay of the singular values or determined based on control performance.